# MindCraft: How Concept Trees Take Shape In Deep Models

### v1

@@ Appendix A The Comparison between Raw Value Matrix and SVD
Second, even after careful tuning, the Raw Value method fails to differentiate semantically distinct concepts, leading to degenerate, flattened tree structures. In Figure 7 (a), the baseline struggles to separate pairs like “diabetes/hypertension” and “2023/2024”, which MindCraft resolves at different hierarchical levels. This failure to capture nuanced distinctions confirms that raw vector space is less informative than the principal axes of transformation identified by SVD.
In summary, this comparison demonstrates that the SVD-based MindCraft method is more robust, sensitive, and efficient. By analyzing representations along principal transformation axes, MindCraft builds a more meaningful and hierarchical understanding of a model’s internal decision-making process.
@@
Figure 7: Compare MindCraft with Raw Value in six scenarios: (a–b) medical diagnosis, (c) daily life, (d) personality evaluation, (e) physics reasoning, and (f) political decision-making. Counterfactual tokens are marked with underlines, and n denotes the number of remaining unbranched concepts.
Appendix B Automated Concept Extraction
@@
Appendix C Experiment Setup
@@
  • TruthfulQA (Lin et al., 2022): A benchmark designed to evaluate whether models generate factually correct and truthful answers to deliberately misleading or adversarial questions. It directly probes concepts such as honesty, reliability, and bias in language generation.
  • AI2 Reasoning Challenge (ARC) (Clark et al., 2018): A collection of grade-school level multiple-choice science questions, split into Easy and Challenge subsets. It tests a model’s ability to apply commonsense knowledge and perform scientific reasoning, beyond simple pattern matching.
  • COPA (Choice of Plausible Alternatives) (Roemmele et al., 2011): A causal reasoning benchmark where the model is given a premise and asked to choose the more plausible cause or effect from two alternatives. It provides a focused evaluation of the model’s ability to capture causal structures in language.
@@
LLM: We employ Qwen2.5-7B-Instruct (Team, 2024), a 7-billion-parameter instruction-tuned large language model developed by Alibaba’s Qwen team as part of the Qwen2.5 series. It adopts a decoder-only Transformer architecture with improvements such as SwiGLU activations and group query attention, and supports very long context lengths of up to 128K tokens. Compared to its predecessor, Qwen2, it demonstrates stronger instruction-following ability, richer knowledge (especially in coding and mathematics), better handling of structured data, and multilingual support across more than 29 languages.
Appendix D The Use of Large Language Models (LLMs)

### v2

@@ B.1 The Comparison between Raw Value Matrix and SVD
Second, even after careful tuning, the Raw Value method is not efficient enough to differentiate semantically distinct concepts, leading to degenerate, flattened tree structures. In Figure 7 (a), the baseline struggles to separate pairs like “diabetes/hypertension” and “2023/2024”, which MindCraft resolves at different hierarchical levels. This inefficiency to capture nuanced distinctions confirms that raw vector space is less informative than the principal axes of transformation identified by SVD.
In summary, this comparison demonstrates that the SVD-based MindCraft method is more robust, sensitive, and efficient. By analyzing representations along principal transformation axes, MindCraft builds a more meaningful and hierarchical understanding of a model’s internal process.
B.2 Ablation Study
B.2.1 Ablation Study on $k$
To investigate the effect of the parameter $k$, which controls the number of principal components retained in the Concept Path, we conduct an ablation study using the same scenario as in Figure 3 (a). As shown in Figure 8, both extremely small and large values of $k$ cannot effectively capture coherent conceptual structures. When $k$ is too small (e.g., $k=1$), the model retains only a single dominant spectral direction, losing semantic distinctions. This indicates that conceptual representations within the model are governed by the collective interaction of multiple dimensions rather than by a few isolated ones, suggesting that concept formation is inherently distributed across several principal components.
Conversely, when $k$ is too large (e.g., $k=100$), excessive noise from less informative components overwhelms the main conceptual signal, resulting in unstable or collapsed Concept Trees. The optimal balance, empirically found at $k=10$, provides a stable and interpretable hierarchy where concept separations emerge at appropriate layers. This demonstrates that moderate values of $k$ are essential for capturing meaningful conceptual dynamics while maintaining robustness.
B.2.2 Ablation Study on $\tau$
We further examine the effect of the separation threshold $\tau$, which determines the layer at which two concepts are considered distinct in the Concept Tree. As illustrated in Figure 9, both overly high and overly low values of $\tau$ fail to yield stable and interpretable structures. When $\tau$ is too high (e.g., $\tau=0.99$), even minor fluctuations in similarity trigger premature branching, causing the model to over-segment representations and produce shallow and flat trees. In contrast, when $\tau$ is too low (e.g., $\tau=0.7$), the model delays concept separation until very late layers, also collapsing distinct semantic branches into overly flat structures. The balanced threshold of $\tau=0.9$ provides the most coherent hierarchy, aligning with the natural emergence of conceptual distinctions observed across layers. This suggests that an appropriate $\tau$ is essential for capturing genuine conceptual divergence without introducing spurious separations.
@@ B.2.2 Ablation Study on $\tau$
Figure 7: Compare MindCraft with Raw Value in six scenarios: (a–b) medical diagnosis, (c) daily life, (d) personality evaluation, (e) physics reasoning, and (f) political decision-making. Counterfactual tokens are marked with underlines, and n denotes the number of unbranched concepts.
(a) $k=1$
(b) $k=10$
(c) $k=50$
(d) $k=100$
Figure 8: Ablation study on $k$, where we use the same scenario as in Figure 3 (a). Neither small nor large values of $k$ can effectively capture meaningful conceptual information.
(a) $\tau=0.99$
(b) $\tau=0.9$
(c) $\tau=0.8$
(d) $\tau=0.7$
Figure 9: Ablation study on $\tau$, where we use the same scenario as in Figure 3 (a). Similar to $k$, both overly high and overly low values of $\tau$ fail to produce stable and meaningful concept separations.
B.2.3 Ablation on LLMs
To further evaluate the generality of the proposed Concept Tree framework, we conducted ablation studies across five representative large language models with varying architectures and parameter scales: Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct and Qwen2.5-32B-Instruct (qwen2025qwen25technicalreport); Mistral-7B-Instruct (jiang2023mistral7b); LLaMA-7B and LLaMA-30B (touvron2023llama)
The results consistently demonstrate that the Concept Tree framework can be effectively applied across diverse model families, confirming its robustness and architectural generality. Despite differences in training objectives and model sizes, all five models exhibit similar hierarchical branching dynamics, indicating that the formation and stabilization of conceptual representations are a shared phenomenon among modern large language models.
(a) Qwen2.5-7B-Instruct
(b) Qwen2.5-14B-Instruct
(c) Qwen2.5-32B-Instruct
(d) Mistral-7B-Instruct
(e) LLaMA-7B
(f) LLaMA-30B
Figure 10: Cross-model comparison across different LLM architectures, including Qwen, Mistral, and LLaMA series, where we use the same scenario as in Figure 3 (a). The results illustrate consistent hierarchical branching dynamics across scales and architectures.
Appendix C Automated Concept Extraction
@@
Appendix D Experiment Setup
@@
  • TruthfulQA (lin2022truthfulqa): A benchmark designed to evaluate whether models generate factually correct and truthful answers to deliberately misleading or adversarial questions. It directly probes concepts such as honesty, reliability, and bias in language generation.
  • AI2 Reasoning Challenge (ARC) (clark2018think): A collection of grade-school level multiple-choice science questions, split into Easy and Challenge subsets. It tests a model’s ability to apply commonsense knowledge and perform scientific reasoning, beyond simple pattern matching.
  • COPA (Choice of Plausible Alternatives) (roemmele2011choice): A causal reasoning benchmark where the model is given a premise and asked to choose the more plausible cause or effect from two alternatives. It provides a focused evaluation of the model’s ability to capture causal structures in language.
@@
LLM: We employ Qwen2.5-7B-Instruct (qwen2025qwen25technicalreport), a 7-billion-parameter instruction-tuned large language model developed by Alibaba’s Qwen team as part of the Qwen2.5 series. It adopts a decoder-only Transformer architecture with improvements such as SwiGLU activations and group query attention, and supports very long context lengths of up to 128K tokens. Compared to its predecessor, Qwen2, it demonstrates stronger instruction-following ability, richer knowledge (especially in coding and mathematics), better handling of structured data, and multilingual support across more than 29 languages.
Appendix E The Use of Large Language Models (LLMs)